OpenStudy (anonymous):

what is an equation of the line in point-slope form that passes through the given point and has the given slope? point: (8,-8) slope 3

OpenStudy (***[isuru]***):

the general format of a straight line is y = mx + b where m = slope and b = intercept since the slope is 3 u can write the equation of the line as y= 3x + b now all u want is to find the value of b... to find b u can use the coordinates (8 , -8) this coordinates means that the line will pass through a point where x = 8 and y = -8 plug these values to the equation y= 3x + b then u will get -8 = (3x8) +b -8 = 24 + b b = -32 that's it!! then u will have the equation of the straight line as y = 3x -32 hope this will help ya:)